# FPGA-based Emulation and Device-Side Management for CXL-based Memory Tiering Systems

### v2

@@ Abstract.
In this study, we first introduce HeteroBox, a configurable emulation platform that leverages real CXL-enabled FPGAs to emulate the performance of various CXL memory architectures. HeteroBox allows one to configure a memory space with multiple regions, each exhibiting distinct CPU-access latency and bandwidth. HeteroBox helps assess the performance of both software-managed and hardware-managed memory tiering systems with high efficiency and fidelity. Based on HeteroBox, we further propose HeteroMem [^1] The code of HeteroBox and HeteroMem will be released after publication., a hardware-managed memory tiering system that operates on the device side. HeteroMem creates an abstraction layer between the CPU and device memory, effectively monitoring data usage and migrating data to faster memory tiers, thus hiding device-side heterogeneity from the CPU. Evaluations with real-world applications show that HeteroMem delivers high performance while keeping heterogeneous memory management fully transparent to the CPU, achieving a 5.1% to 16.2% performance improvement over existing memory tiering solutions.
* Co-first authors.$\dagger$ Corresponding author.
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 52nd IEEE/ACM International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japanisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
The increasing demand for DRAM in modern data centers has led to its significant share in both operational costs and power consumption, consuming approximately 37% of the total cost and 33% of the power, according to Meta (Maruf et al., 2023). This dependency on DRAM is further entrenched by the DDR standard, which is designed around DRAM-specific access and maintenance commands, thereby limiting the integration of emerging memory technologies.
The emerging Compute Express Link (CXL) technique enables CPUs to access a variety of memory types by offering unified load-store semantics. This breakthrough paves the way for the adoption of alternative memory technologies, such as Managed DRAM (Lee et al., 2019), ReRAM (Chen, 2020), and 3D-XPoint/Optane (Hady et al., 2017), which promise larger capacities and lower costs (Dulloor et al., 2016) (Eisenman et al., 2018) (Kassa et al., 2021), potentially reducing the Total Cost of Ownership (TCO) for data centers (Sharma et al., 2024). As shown in Figure 1-(a), different types of memory media can be integrated into single or multiple CXL memory devices. CPU can directly access the expanded memory via CXL channels, avoiding modifying the DDR memory controllers in existing CPUs. In addition to enhancing the diversity of memory media connected to the CPU, the CXL standard supports connecting the host to multiple endpoint devices via switches, enabling larger-scale memory expansion. Additionally, CXL 3.0 introduces the capability for cascaded CXL switches (foundation, [n. d.]). As illustrated in Figure 1-(b), future CXL memory systems can feature a CXL switch connected to a host, another CXL switch, or a CXL memory device, thereby creating a complex topology.
@@ 1. Introduction
Figure 1. CXL-extended Heterogeneous Memory System
Prior research has endeavored to refine systems equipped with CXL memory from both system (Maruf et al., 2023; Lee et al., 2023) and architectural perspectives (Zhou et al., 2024; Zhong et al., 2024). The main idea of existing solutions is to leverage the access locality of workloads, identifying frequently accessed hot data and placing it into fast memory tiers. The remaining cold data is migrated to slow memory tiers to make room for hot data (Ren et al., 2023; Sha et al., 2023; Dong et al., 2010a). However, prior works face challenges in optimizing the performance of the emerging CXL-extended heterogeneous memory systems. The reasons are as follows:
@@
Challenge 2: Low Performance of Software-Managed Memory Tiering System - As detailed in Section 2, previous software-based memory tiering systems, which rely on various memory access profiling mechanisms including PTE scanning, hint-fault monitoring, and PMU sampling, struggle to accurately and efficiently identify hot and cold data due to hardware limitations. Meanwhile, the overhead of software-based data migration is significant.
@@
  • We propose HeteroBox, a configurable emulation platform leveraging real CXL-enabled FPGAs to emulate the performance of various CXL memory architectures. By carefully designing its architecture, we enable HeteroBox to support multiple emulated regions, each with distinct latency and bandwidth characteristics (Section 3).
  • We propose HeteroMem, a hardware-managed memory tiering system operating on the device side. We carefully design the architecture of HeteroMem to accurately and efficiently detect hot and cold data, and to manage data placement through effective data migration (Section 4).
  • We evaluate HeteroMem’s performance on the HeteroBox emulation platform and explore various design choices for hardware-managed memory tiering systems. Our findings demonstrate that HeteroMem achieves a performance improvement of 5.7% $\sim$ 17.6% over several existing memory tiering systems, highlighting its effectiveness and potential as a superior memory management solution (Section 6).
2. Background & Motivation
Table 1. Comparison of HeteroMem with other Memory Tiering Systems
@@ 2. Background & Motivation
AutoNUMA (Corbet, 2012) Hint-fault monitoring High Low Moved by CPU High Low No Yes
TPP (Maruf et al., 2023) Hint-fault monitoring High Low Moved by CPU High Low No Yes
DAMON (Park, 2021) PTE Scanning High Low Moved by CPU High Low No Yes
MEMTIS (Lee et al., 2023) PMU Sampling Median Median Moved by CPU High Low No Yes
HeMem (Raybuck et al., 2021) PMU Sampling Median Median Moved by CPU High Low No Yes
MemPod (Prodromou et al., 2017) Hardware Counters Low Median Moved by Hardware Low High Yes No
Neomem (Zhou et al., 2024) Hardware Hot Page Profiler Low High(only for hot pages) Moved by CPU High Low No Yes
@@
2.1. CXL-based Memory Expansion
@@
The Compute Express Link (CXL) protocol(foundation, [n. d.]), based on the PCIe 5.0 physical layer, facilitates a cache-coherent, byte-addressable interconnect through high-efficiency SerDes links. This protocol composes three primary sub-protocols: CXL.io, CXL.cache, and CXL.mem. The CXL.mem sub-protocol is particularly noteworthy as it allows for direct CPU access to memory devices at cache-line granularity through the CXL interface. Unlike traditional DDR memory architectures with memory controllers in the host CPU, CXL incorporates memory controllers on the device side. This decouples the architecture, allowing flexible integration of various memory technologies into CXL memory controllers while preserving the host CPU design. Such flexibility is crucial for meeting specific capacity, performance, and cost-effectiveness needs of modern memory systems (Hynix, [n. d.]; Sumsung, [n. d.]; Li et al., 2023a; Boles et al., 2023).
@@
CXL 2.0 introduces the CXL switch, which supports a single level of switching. The CXL switch allows one or multiple hosts to connect to one or more CXL devices. CXL 3.0 allows multiple switch levels, which enables more complex CXL memory device topologies. CXL 3.0 uses Port Based Routing (PBR) for inter-switch links to route CXL messages, allowing the software running on the host to perceive a flat topology ending at the leaf switch port. However, using a CXL switch introduces additional latency. According to (Li et al., 2023a), each CXL switch that a CXL request passes through adds about 70ns of latency. In a complex CXL memory system consisting of multiple switch levels, memory requests to different devices may pass through varying numbers of switches, resulting in a memory system with heterogeneous latency and bandwidth characteristics to the host CPU.
@@
However, due to the limited availability of real-world CXL-enabled CPUs and devices, the evaluation of CXL-extended memory systems faces significant constraints. Specifically, previous work often relies on remote NUMA node emulation (Maruf et al., 2023; Lee et al., 2023) or software-based simulation methods (Jang et al., 2023) to assess the performance of CXL-based heterogeneous memory systems. But real CXL-extended memory exhibits different latency and bandwidth characteristics compared to remote NUMA nodes according to (Sun et al., 2023). Furthermore, CXL decouples the CPU and memory interface, allowing for more complex modifications on the device side, which cannot be modeled by NUMA emulation-based platforms. Software-based simulation methods, however, struggle with handling large workloads and assessing server CPU performance at a real scale due to low simulation speeds. For instance, the simulation speed of gem5 (Lowe-Power et al., 2020) in full system mode is approximately 100 kilo instructions per second (KIPS), which is tens of thousands of times slower than emulation methods. While some studies utilize CXL-enabled CPUs and devices (Sun et al., 2023; Zhang et al., 2023; Sano et al., 2023) for evaluation, their configurations are restricted by the available hardware and can hardly adapt to other settings. These limitations highlight the need for an emulation platform that can efficiently and accurately evaluate the performance of CXL-extended heterogeneous memory systems.
@@
2.2. Memory Tiering System
Many studies have focused on developing memory tiering systems to enhance the performance of heterogeneous memory systems. Specifically, we examine three critical aspects of memory tiering system design: (1) Memory Access Profiling, (2) Data Movement, and (3) Flexibility and Compatibility. Table 1 provides a summary comparison of prior work.
Figure 2. Overhead of Software Profiling and Data Movement
Memory Access Profiling: The memory access profiling mechanism is crucial for managing heterogeneous memory systems and achieving performance improvements. However, compared to hardware-based methods, software-based memory access profiling methods suffer from high overhead and low resolution. Software-based methods primarily consist of three types: hint-fault monitoring, PTE scanning, and PMU-based sampling. Hint-fault monitoring periodically poisons a portion of the memory pages, invalidating their PTEs and causing page faults when they are next accessed. TPP (Maruf et al., 2023) and AutoNUMA (Corbet, 2012) utilize hint-fault monitoring to detect hot data in slow memory. PTE scanning uses the access bit in the PTE to detect hot data. When a page is accessed, the MMU automatically sets the access bit. A kernel thread periodically scans the memory space to check whether a page has been accessed during the last period. AMP (Heo et al., 2022) uses PTE scanning to detect hot data in slow memory, while TMTS (Duraisamy et al., 2023) uses PTE scanning to detect cold data in fast memory. PMU sampling leverages specific hardware counters in the CPU to profile hot data in slow memory. For example, Intel CPUs support using PEBS counters to record LLC miss events. Each time an LLC miss event occurs, the PEBS counter increments by one. When the counter reaches a user-defined threshold, it overflows and records the access address. HeMem (Raybuck et al., 2021) and MEMTIS (Lee et al., 2023) use PMU sampling to detect hot data in slow memory.
However, as discussed in previous work (Zhou et al., 2024), these mechanisms fail to achieve both low overhead and high resolution simultaneously. As shown in Figure 2-(a), we use DAMON (Park, 2021) to profile the CPU overhead introduced by PTE scanning as an example. DAMON is a data access monitoring framework subsystem for the Linux kernel, which can be configured to profile memory access with different spatial and temporal granularities. We vary the space granularity from 50 regions to 1050 regions and the time granularity from scanning the page table every 50ms to every 1050ms. We visualize the overhead as a heat map, where lighter colors indicate heavier overhead. We observe significant overhead when using small spatial or temporal granularities, with a maximum cost of 88.15% CPU time.
@@ 2.2. Memory Tiering System
Figure 3. Overview of HeteroBox Emulation Platform.
Data Movement: After hot data is detected in slow memory, a data movement mechanism is required to migrate it to fast memory. Typically, the CPU can migrate data between fast and slow memory. Many previous works utilize the migrate_pages API in the Linux kernel for this purpose. However, as demonstrated in Nimble (Yan et al., 2019), using the CPU for data migration incurs significant overhead and reduced bandwidth due to the software stack in the OS, including tasks like checking permissions and unmapping and remapping pages. Although some works attempt to mitigate these limitations by using transactional or parallel migration, they still rely on the CPU for data migration, which inevitably causes CPU overhead. Furthermore, using the CPU to migrate data between CXL memory devices requires reading the data out and then writing it back, which consumes the limited bandwidth of the CXL link. As shown in Figure 2-(b), we compare the bandwidth of using custom hardware logic, as described in Section 4, to migrate data versus using the CPU to migrate data. For both methods, we use the 4KB page granularity and record the maximum number of pages that can be migrated per second. We observe that using custom hardware logic to migrate data achieves 12.9x higher bandwidth compared to using the CPU.
@@
Other works employ hardware-based memory management for better performance. For example, MemPod (Prodromou et al., 2017) modifies the CPU-side memory controller to group different types of memory into several pods. The MemPod logic in the memory controller optimizes the performance of the heterogeneous memory system without CPU involvement. These hardware-based heterogeneous memory management systems remain transparent to CPU. However, due to architectural limitations, they require modifications to CPU hardware, making them unsuitable to adapt to various configurations of CXL memory.
A recent state-of-the-art work, NeoMem (Zhou et al., 2024), proposes integrating a hardware unit within the CXL memory device, which communicates with the host CPU to provide memory access information. However, NeoMem relies on CPU to process this information and perform subsequent data movement, making it non-transparent to CPU and reducing overall flexibility and compatibility.
@@
3. HeteroBox Platform
3.1. Overview of HeteroBox Emulation Platform
@@
3.2. Emulation Logic Design
@@ 3.3. Configuration Process
4. HeteroMem Solution
4.1. Overview of HeteroMem System
Figure 4. System Overview of HeteroMem
@@ 4.1. Overview of HeteroMem System
Figure 5. Remapping Unit Working Process.
Figure 6. Block Diagram of the Remapping Unit
@@ 4.2. Remapping Unit
Figure 7. Profiling Unit Overview.
@@ 4.3. Profiling Unit
The Profiling Unit is responsible for profiling the hotness and coldness of data in memory. In our HeteroMem design, we profile data hotness and coldness at a 4kB page granularity. The Profiling Unit is linked after the Remapping Unit, off the critical path, to capture the dPA trace. We profile only read requests, as write requests are generally not on the critical path of program execution. As shown in Figure 7-(a), in the case of a CXL-extended memory system with two tiers, the Profiling Unit splits the read requests into two streams: one targeting fast memory and the other targeting slow memory. For the read request stream targeting fast memory, the Profiling Unit detects cold pages based on their access history. These detected cold pages are temporarily buffered in a cold pages buffer. Since cold pages tend to remain cold for a relatively long time (Duraisamy et al., 2023), they typically remain cold when fetched from the buffer in subsequent processes. For the read request stream targeting slow memory, the Profiling Unit measures the hotness of each page based on its access frequency, reporting a page as hot if its hotness exceeds a given threshold. Upon detecting a hot page, the Profiling Unit fetches a cold page from the cold pages buffer and sends the detected hot page and the fetched cold page to the Remapping Unit. The Remapping Unit then manages the data movement, swapping pages between fast memory and slow memory. The number of these hot and cold page pairs sent to the Remapping Unit is limited by a given constraint (e.g., 32 pairs of pages per 100,000 cycles).
[…]
@@
Coldness Profiling: As detailed in Section 4.D, the Migration Unit uses a swap mechanism to move hot data to fast memory. Therefore, when a hot page is detected in slow memory, we need to find a cold page in fast memory to swap its location with the detected hot page. We use a ping-pong bitmap to record the access history of each page in fast memory. The ping-pong bitmap structure consists of two bitmap arrays, each entry being a bit corresponding to a page in fast memory. In a given period, one bitmap, say bitmap $A$, is used to record memory access in the current period. When Profiling Unit receives a memory request, it sets the bit corresponding to the accessed page in bitmap $A$. Meanwhile, Profiling Unit scans the other bitmap, say bitmap $B$, to identify pages with their corresponding bit in bitmap $B$ unset as cold pages. In the next period, we reset all bits in bitmap $B$ and switch the functions of bitmap $A$ and bitmap $B$. This approach ensures the detected cold pages were not accessed during the last period, thus accurately identifying cold pages.
4.4. Migration Unit

### v3

# Exploring Memory Tiering Systems in the CXL Era via FPGA-based Emulation and Device-Side Management

@@ Abstract
In this study, we first introduce HeteroBox, a configurable emulation platform that leverages real CXL-enabled FPGAs to emulate the performance of various CXL memory architectures. HeteroBox allows one to configure a memory space with multiple regions, each exhibiting distinct CPU-access latency and bandwidth. HeteroBox helps assess the performance of both software-managed and hardware-managed memory tiering systems with high efficiency and fidelity. Based on HeteroBox, we further propose HeteroMem, a hardware-managed memory tiering system that operates on the device side. HeteroMem creates an abstraction layer between the CPU and device memory, effectively monitoring data usage and migrating data to faster memory tiers, thus hiding device-side heterogeneity from the CPU. Evaluations with real-world applications show that HeteroMem delivers high performance while keeping heterogeneous memory management fully transparent to the CPU, achieving a 5.1% to 16.2% performance improvement over existing memory tiering solutions.
I Introduction
The increasing demand for DRAM in modern data centers has led to its significant share in both operational costs and power consumption, consuming approximately 37% of the total cost and 33% of the power, according to Meta [40]. This dependency on DRAM is further entrenched by the DDR standard, which is designed around DRAM-specific access and maintenance commands, thereby limiting the integration of emerging memory technologies.
[…]
Figure 1: CXL-extended Heterogeneous Memory System
The emerging Compute Express Link (CXL) technique enables CPUs to access a variety of memory types by offering unified load-store semantics. This breakthrough paves the way for the adoption of alternative memory technologies, such as Managed DRAM [32], ReRAM [10], and 3D-XPoint/Optane [23], which promise larger capacities and lower costs [18] [20] [28], potentially reducing the Total Cost of Ownership (TCO) for data centers [50]. As shown in Figure 1-(a), different types of memory media can be integrated into single or multiple CXL memory devices. CPU can directly access the expanded memory via CXL channels, avoiding modifying the DDR memory controllers in existing CPUs. In addition to enhancing the diversity of memory media connected to the CPU, the CXL standard supports connecting the host to multiple endpoint devices via switches, enabling larger-scale memory expansion. Additionally, CXL 3.0 introduces the capability for cascaded CXL switches [21]. As illustrated in Figure 1-(b), future CXL memory systems can feature a CXL switch connected to a host, another CXL switch, or a CXL memory device, thereby creating a complex topology.
@@ I Introduction
Prior research has endeavored to refine systems equipped with CXL memory from both system [40, 34] and architectural perspectives [61, 60]. The main idea of existing solutions is to leverage the access locality of workloads, identifying frequently accessed hot data and placing it into fast memory tiers. The remaining cold data is migrated to slow memory tiers to make room for hot data [45, 49, 16]. However, prior works face challenges in optimizing the performance of the emerging CXL-extended heterogeneous memory systems. The reasons are as follows:
@@
Challenge 2: Low Performance of Software-Managed Memory Tiering System - As detailed in Section II, previous software-based memory tiering systems, which rely on various memory access profiling mechanisms including PTE scanning, hint-fault monitoring, and PMU sampling, struggle to accurately and efficiently identify hot and cold data due to hardware limitations. Meanwhile, the overhead of software-based data migration is significant.
@@
  • We propose HeteroBox, a configurable emulation platform leveraging real CXL-enabled FPGAs to emulate the performance of various CXL memory architectures. By carefully designing its architecture, we enable HeteroBox to support multiple emulated regions, each with distinct latency and bandwidth characteristics (Section III).
  • We propose HeteroMem, a hardware-managed memory tiering system operating on the device side. We carefully design the architecture of HeteroMem to accurately and efficiently detect hot and cold data, and to manage data placement through effective data migration (Section IV).
  • We evaluate HeteroMem’s performance on the HeteroBox emulation platform and explore various design choices for hardware-managed memory tiering systems. Our findings demonstrate that HeteroMem achieves a performance improvement of 5.7% $\sim$ 17.6% over several existing memory tiering systems, highlighting its effectiveness and potential as a superior memory management solution (Section VI).
II Background & Motivation
TABLE I: Comparison of HeteroMem with other Memory Tiering Systems
@@ II Background & Motivation
AutoNUMA [14] Hint-fault monitoring High Low Moved by CPU High Low No Yes
TPP [40] Hint-fault monitoring High Low Moved by CPU High Low No Yes
DAMON [41] PTE Scanning High Low Moved by CPU High Low No Yes
MEMTIS [34] PMU Sampling Median Median Moved by CPU High Low No Yes
HeMem [43] PMU Sampling Median Median Moved by CPU High Low No Yes
MemPod [42] Hardware Counters Low Median Moved by Hardware Low High Yes No
Neomem [61] Hardware Hot Page Profiler Low High(only for hot pages) Moved by CPU High Low No Yes
@@
II-A CXL-based Memory Expansion
@@
The Compute Express Link (CXL) protocol[21], based on the PCIe 5.0 physical layer, facilitates a cache-coherent, byte-addressable interconnect through high-efficiency SerDes links. This protocol composes three primary sub-protocols: CXL.io, CXL.cache, and CXL.mem. The CXL.mem sub-protocol is particularly noteworthy as it allows for direct CPU access to memory devices at cache-line granularity through the CXL interface. Unlike traditional DDR memory architectures with memory controllers in the host CPU, CXL incorporates memory controllers on the device side. This decouples the architecture, allowing flexible integration of various memory technologies into CXL memory controllers while preserving the host CPU design. Such flexibility is crucial for meeting specific capacity, performance, and cost-effectiveness needs of modern memory systems [26, 51, 36, 9].
@@
CXL 2.0 introduces the CXL switch, which supports a single level of switching. The CXL switch allows one or multiple hosts to connect to one or more CXL devices. CXL 3.0 allows multiple switch levels, which enables more complex CXL memory device topologies. CXL 3.0 uses Port Based Routing (PBR) for inter-switch links to route CXL messages, allowing the software running on the host to perceive a flat topology ending at the leaf switch port. However, using a CXL switch introduces additional latency. According to [36], each CXL switch that a CXL request passes through adds about 70ns of latency. In a complex CXL memory system consisting of multiple switch levels, memory requests to different devices may pass through varying numbers of switches, resulting in a memory system with heterogeneous latency and bandwidth characteristics to the host CPU.
@@
However, due to the limited availability of real-world CXL-enabled CPUs and devices, the evaluation of CXL-extended memory systems faces significant constraints. Specifically, previous work often relies on remote NUMA node emulation [40, 34] or software-based simulation methods [27] to assess the performance of CXL-based heterogeneous memory systems. But real CXL-extended memory exhibits different latency and bandwidth characteristics compared to remote NUMA nodes according to [52]. Furthermore, CXL decouples the CPU and memory interface, allowing for more complex modifications on the device side, which cannot be modeled by NUMA emulation-based platforms. Software-based simulation methods, however, struggle with handling large workloads and assessing server CPU performance at a real scale due to low simulation speeds. For instance, the simulation speed of gem5 [38] in full system mode is approximately 100 kilo instructions per second (KIPS), which is tens of thousands of times slower than emulation methods. While some studies utilize CXL-enabled CPUs and devices [52, 59, 48] for evaluation, their configurations are restricted by the available hardware and can hardly adapt to other settings. These limitations highlight the need for an emulation platform that can efficiently and accurately evaluate the performance of CXL-extended heterogeneous memory systems.
@@
II-B Memory Tiering System
Many studies have focused on developing memory tiering systems to enhance the performance of heterogeneous memory systems. Specifically, we examine three critical aspects of memory tiering system design: (1) Memory Access Profiling, (2) Data Movement, and (3) Flexibility and Compatibility. Table I provides a summary comparison of prior work.
Figure 2: Overhead of Software Profiling and Data Movement
Memory Access Profiling: The memory access profiling mechanism is crucial for managing heterogeneous memory systems and achieving performance improvements. However, compared to hardware-based methods, software-based memory access profiling methods suffer from high overhead and low resolution. Software-based methods primarily consist of three types: hint-fault monitoring, PTE scanning, and PMU-based sampling. Hint-fault monitoring periodically poisons a portion of the memory pages, invalidating their PTEs and causing page faults when they are next accessed. TPP [40] and AutoNUMA [14] utilize hint-fault monitoring to detect hot data in slow memory. PTE scanning uses the access bit in the PTE to detect hot data. When a page is accessed, the MMU automatically sets the access bit. A kernel thread periodically scans the memory space to check whether a page has been accessed during the last period. AMP [24] uses PTE scanning to detect hot data in slow memory, while TMTS [19] uses PTE scanning to detect cold data in fast memory. PMU sampling leverages specific hardware counters in the CPU to profile hot data in slow memory. For example, Intel CPUs support using PEBS counters to record LLC miss events. Each time an LLC miss event occurs, the PEBS counter increments by one. When the counter reaches a user-defined threshold, it overflows and records the access address. HeMem [43] and MEMTIS [34] use PMU sampling to detect hot data in slow memory.
However, as discussed in previous work [61], these mechanisms fail to achieve both low overhead and high resolution simultaneously. As shown in Figure 2-(a), we use DAMON [41] to profile the CPU overhead introduced by PTE scanning as an example. DAMON is a data access monitoring framework subsystem for the Linux kernel, which can be configured to profile memory access with different spatial and temporal granularities. We vary the space granularity from 50 regions to 1050 regions and the time granularity from scanning the page table every 50ms to every 1050ms. We visualize the overhead as a heat map, where lighter colors indicate heavier overhead. We observe significant overhead when using small spatial or temporal granularities, with a maximum cost of 88.15% CPU time.
@@ II-B Memory Tiering System
Figure 3: Overview of HeteroBox Emulation Platform.
Data Movement: After hot data is detected in slow memory, a data movement mechanism is required to migrate it to fast memory. Typically, the CPU can migrate data between fast and slow memory. Many previous works utilize the migrate_pages API in the Linux kernel for this purpose. However, as demonstrated in Nimble [58], using the CPU for data migration incurs significant overhead and reduced bandwidth due to the software stack in the OS, including tasks like checking permissions and unmapping and remapping pages. Although some works attempt to mitigate these limitations by using transactional or parallel migration, they still rely on the CPU for data migration, which inevitably causes CPU overhead. Furthermore, using the CPU to migrate data between CXL memory devices requires reading the data out and then writing it back, which consumes the limited bandwidth of the CXL link. As shown in Figure 2-(b), we compare the bandwidth of using custom hardware logic, as described in Section IV, to migrate data versus using the CPU to migrate data. For both methods, we use the 4KB page granularity and record the maximum number of pages that can be migrated per second. We observe that using custom hardware logic to migrate data achieves 12.9x higher bandwidth compared to using the CPU.
@@
Other works employ hardware-based memory management for better performance. For example, MemPod [42] modifies the CPU-side memory controller to group different types of memory into several pods. The MemPod logic in the memory controller optimizes the performance of the heterogeneous memory system without CPU involvement. These hardware-based heterogeneous memory management systems remain transparent to CPU. However, due to architectural limitations, they require modifications to CPU hardware, making them unsuitable to adapt to various configurations of CXL memory.
A recent state-of-the-art work, NeoMem [61], proposes integrating a hardware unit within the CXL memory device, which communicates with the host CPU to provide memory access information. However, NeoMem relies on CPU to process this information and perform subsequent data movement, making it non-transparent to CPU and reducing overall flexibility and compatibility.
@@
III HeteroBox Platform
III-A Overview of HeteroBox Emulation Platform
@@
III-B Emulation Logic Design
@@ III-C Configuration Process
IV HeteroMem Solution
IV-A Overview of HeteroMem System
Figure 4: System Overview of HeteroMem
@@ IV-A Overview of HeteroMem System
Figure 5: Remapping Unit Working Process.
Figure 6: Block Diagram of the Remapping Unit
@@ IV-B Remapping Unit
Figure 7: Profiling Unit Overview.
@@ IV-C Profiling Unit
The Profiling Unit is responsible for profiling the hotness and coldness of data in memory. In our HeteroMem design, we profile data hotness and coldness at a 4kB page granularity. The Profiling Unit is linked after the Remapping Unit, off the critical path, to capture the dPA trace. We profile only read requests, as write requests are generally not on the critical path of program execution. As shown in Figure 7-(a), in the case of a CXL-extended memory system with two tiers, the Profiling Unit splits the read requests into two streams: one targeting fast memory and the other targeting slow memory. For the read request stream targeting fast memory, the Profiling Unit detects cold pages based on their access history. These detected cold pages are temporarily buffered in a cold pages buffer. Since cold pages tend to remain cold for a relatively long time [19], they typically remain cold when fetched from the buffer in subsequent processes. For the read request stream targeting slow memory, the Profiling Unit measures the hotness of each page based on its access frequency, reporting a page as hot if its hotness exceeds a given threshold. Upon detecting a hot page, the Profiling Unit fetches a cold page from the cold pages buffer and sends the detected hot page and the fetched cold page to the Remapping Unit. The Remapping Unit then manages the data movement, swapping pages between fast memory and slow memory. The number of these hot and cold page pairs sent to the Remapping Unit is limited by a given constraint (e.g., 32 pairs of pages per 100,000 cycles).
Hotness Profiling: For memory read requests targeting fast memory, we use a Count-Min Sketch [15] structure to profile the hotness of each page. Count-Min Sketch is a hash-based algorithm that can estimate heavy hitters within a stream. As shown in Figure 7-(b), the main body of the Count-Min Sketch is a $\textbf{D}*\textbf{W}$ counter array, where D is the depth of the array and W is the width. The W entries in the same row are referred to as a lane. When a request comes in, its dPA is mapped to D different lanes using D different hash functions. The mapped counters are incremented by one, and when the counters overflow, they remain at their maximum value. The minimum value of the D mapped counters is then sent to the subsequent hot page detection logic. If this minimum value exceeds a given threshold, the accessed page is marked as hot. Each counter has an associated hot bit that is set when the corresponding page is detected as hot. A page will not be reported as hot if all of its D hot bits are already set, thus avoiding repeated reporting of the same hot page. All counters are reset periodically. Since the counters in the Count-Min Sketch provide an approximate value of access frequency and the error bound of the Count-Min Sketch is determined by the length of the stream, we control the reset period to limit the sketch’s error bound.
@@
Coldness Profiling: As detailed in Section IV.D, the Migration Unit uses a swap mechanism to move hot data to fast memory. Therefore, when a hot page is detected in slow memory, we need to find a cold page in fast memory to swap its location with the detected hot page. We use a ping-pong bitmap to record the access history of each page in fast memory. The ping-pong bitmap structure consists of two bitmap arrays, each entry being a bit corresponding to a page in fast memory. In a given period, one bitmap, say bitmap $A$, is used to record memory access in the current period. When Profiling Unit receives a memory request, it sets the bit corresponding to the accessed page in bitmap $A$. Meanwhile, Profiling Unit scans the other bitmap, say bitmap $B$, to identify pages with their corresponding bit in bitmap $B$ unset as cold pages. In the next period, we reset all bits in bitmap $B$ and switch the functions of bitmap $A$ and bitmap $B$. This approach ensures the detected cold pages were not accessed during the last period, thus accurately identifying cold pages.
IV-D Migration Unit